A solid consists of a mixture of NaNO3 and Mg(NO3)2. When 6.50 g of the solid is dissolved in 50.0 g water, the freezing point is lowered by 5.50°C. What is the composition, by mass, of the solid?
Mg(NO3)2
?g
NaNO3
?g
Answer #1

Step1 Let mass of NaNO3 = x g; Mass of Mg(NO3)2 = [6.5 -x] g
Step2 Moles of NaNO3 = (x/85) ; Moles of Mg(NO3)2 =[6.5-x]/148 g
Step3 Molality of NaNO3 =(x/85) x(1000/50) = 20x/85 = 4x/17
Step4 Molality of Mg(NO3)2 = [6.5 -x]/148 x [1000/50] = 20[6.5 -x]/148
Step5 Δ Tf = Kf x m

Step6 Combined molality = [4x/17] + 20[6.5 -x]/148

Step7      [4x/17]* (1.86) + 20[6.5-x]* (1.86)/148 = 5.5   ; solve for x